OpenStudy (anonymous):

write the standar form of the equation of the equation of the circle with the radius and the center r=1 (h,k) (0,1)

OpenStudy (anonymous):

whats x and y?

OpenStudy (anonymous):

You leave those as x and y. just plug in what you have. and i made a type it should be a + between them (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

you have h,k, and r, simply plug them in (x-0)^2+(y-1)^2=1

OpenStudy (anonymous):

general form of the equation now?

OpenStudy (anonymous):

general form?